# RNBF: Real-Time RGB-D Based Neural Barrier Functions for Safe Robotic Navigation

### v1

@@ Abstract
Autonomous safe navigation in unstructured and novel environments poses significant challenges, especially when environment information can only be provided through low-cost vision sensors. Although safe reactive approaches have been proposed to ensure robot safety in complex environments, many base their theory off the assumption that the robot has prior knowledge on obstacle locations and geometries. In this paper, we present a real-time, vision-based framework that constructs continuous, first-order differentiable Signed Distance Fields (SDFs) of unknown environments without any pre-training. Our proposed method ensures full compatibility with established SDF-based reactive controllers. To achieve robust performance under practical sensing conditions, our approach explicitly accounts for noise in affordable RGB-D cameras, refining the neural SDF representation online for smoother geometry and stable gradient estimates. We validate the proposed method in simulation and real-world experiments using a Fetch robot.
@@ I INTRODUCTION
Safety in robotics is a critical challenge, especially when operating in unstructured environments where obstacle states and geometries are unavailable [1]. In recent years, Control Barrier Functions (CBFs) were widely studied due to their ability to provide safety guarantees for any control affine systems and adopted in various applications including vehicle cruise control [2], indoor navigation [3], quadrotor control [4], legged robot locomotion [5], and etc. All the aforementioned works assume that the robot has prior knowledge of obstacle states and geometries, or even a predefined map. However, for robot navigation in unknown environments, such assumptions are unrealistic and the robot can only access noisy environment information from its onboard sensors.
CBFs’s performance in enforcing robot safety relies not only on the robot’s ability to collect accurate nearby obstacle state information but also on how well the collected information is processed to meet the requirement of valid CBFs. Therefore, when environment descriptions are unavailable or noisy, the safety guarantee of robot navigation is jeopardized. Building valid CBFs adaptively using noisy data collected from robots’ on-board sensors in real time remains a challenge. Many recent works use LiDAR sensors to construct CBFs and ensure robot safety in unknown environments [6, 7, 8, 9]. However, LiDAR systems can be expensive and may not always be available in practical deployments. Instead, we focus on leveraging more affordable and accessible sensing modalities, such as RGB-D cameras.
Seldom works have investigated perception-based CBFs using RGB-D camera inputs. Existing vision-based CBF pipelines utilize machine-learning techniques to incorporate robot observations into CBF formulations. For example, [10] leverages a conditional Generative Adversarial Network to approximate CBFs given RGB-D camera inputs; [11] predicts future RGB-D observations with Neural Radiance Fields (NeRFs) and feeds them into discrete-time CBFs. These methods require significant offline training and their performance can be affected by the quality of their training dataset. Additionally, using CBFs in discrete-time settings may weaken their safety guarantees when the update frequency is not fast enough. Other works, like [12], approximate mathematically CBFs and their gradients from point cloud data collected through LiDARs or depth cameras.
Among the methods listed above, it is noteworthy that none has explicitly considered the effects of sensory noises on perception-based CBF. Further, most existing vision-based CBFs have made approximations to standard CBFs to accommodate robot observation formats. Assumptions that CBFs are founded on, such as the continuity and first-order differentiability of the barrier function are typically relaxed, weakening their safety guarantees. Additionally, altering standard CBFs in vision-based safe navigation hinders the adaptation of well-developed none-sensor-based safe reactive control algorithms into vision-based robot platforms. Many safe reactive controllers, including CBFs, are developed assuming the system’s access to signed distance fields (SDFs) and their gradients of the obstacles near the robot [13],[2],[14],[15], [16] and [3]. In robot safe navigation, SDFs refer to the orthogonal distance of a given point to the boundary of an unsafe set.
@@
Therefore, in this paper, we offer an online barrier function generation framework capable of smoothing sensor noises and constructing continuous and first-order differentiable neural SDFs in real-time given RGB-D camera inputs without pre-training for any unknown environments. With the advancements presented by NeRF[17], the quality and efficiency of environmental reconstruction have significantly improved. However, limited work explored generating SDFs using NeRF, most of which were designed for an offline setting [18, 19]. A notable exception is iSDF[20], which introduced a method for real-time SDF generation. Our vision pipeline builds on this foundation to generate real-time, continuous, and differentiable SDFs, specifically designed for constructing continuous CBFs for safe robot navigation. Instead of altering CBFs to accommodate vision observations, as in prior works, we establish continuous and first-order differentiable SDFs online, which enables our methods to be readily adaptable to SDF-based safe reactive controllers such as standard CBF-QPs [21] and Dynamical System Based Modulations [15]. Our approach is validated in both gazebo simulations and real-life hardware experiments using the Fetch robot in static and quasi-static obstacle environments. Our key contributions are as follows:
  • We introduce an approach to generate SDFs from raw RGB-D data using an online neural representation.
  • To account for depth noise inherent in affordable sensors like RGB-D cameras, we mitigate its impact on neural SDF representations adopting the Huber loss [22].
@@ II Problem Formulation
Given a continuously differentiable function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $h$ is a barrier function if the safe set $C$ (outside the obstacle), the boundary set $\partial C$ (on the boundary of the obstacle), and the unsafe set $\neg C$ (inside the obstacles) of the system can be defined as in (2), (3), (4[21].
@@ III-A Control Barrier Function
$C\text{ is set invariant}\iff\dot{h}(x)\geq 0\;\forall x\in\partial C\$ (5)
Control Barrier Functions are designed by extending Nagumo set invariance theorem to a “control” version[21], where the condition $\forall x\in\partial C$ is rewritten mathematically using an extended $\mathcal{K}_{\infty}$ function $\alpha$,
@@
An extended $\mathcal{K}_{\infty}$ function is a function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ that is strictly increasing and with $\alpha(0)=0$ ; that is, extended $\mathcal{K}_{\infty}$ functions are defined on the entire real line (,)($\infty$,$\infty$)( - ∞ , ∞ ).
@@
$u_{\text{cbf}}=\operatorname*{arg\,min}_{{u}\in\mathbb{R}^{m}}\frac{1}{2}||u-u_{\text{nom}}||_{2}^{2}$ (7)
$L_{f}h(x)+L_{g}h(x)u\geq-\alpha(h(x)).$ (8)
@@ III-B Signed Distance Function (SDF)
When representing three-dimensional geometry, a signed distance field $h_{\text{sdf}}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ is a scalar field that maps a 3D coordinate $\xi=[p_{x},p_{y},p_{z}]^{\top}$ to the scalar signed distance value [20]. The surface $S$ is the zero level-set of the field:
@@ IV-A1 Core Methodology
Our real-time SDF generation pipeline (Figure 3, “SDF Generation” block) is inspired by iSDF [20]. Firstly, Our pipeline utilizes a sequence of posed segmented depth images obtained from the segmentation module. Following established practices in RGB-D scene understanding [23], segmenting out the floor proved crucial for developing an efficient barrier function. Yet, the segmentation module can be replaced by any modern segmentation algorithm, for instance, segment anything [24]. For simplification, we opted for a computationally efficient colour-based segmentation in the RGB image to identify the floor region, subsequently masking the corresponding pixels in the depth image. At the heart of our pipeline lies a Multilayer Perceptron (MLP), which transforms a three-dimensional point $\xi$ into its corresponding signed distance. Starting with a randomly weighted model, we dynamically refine the MLP in real-time, leveraging incoming depth measurements. During each update cycle, we strategically select a subset of frames from the depth image stream, employing active sampling to prevent catastrophic forgetting. For every chosen frame, a batch of points is generated by sampling random pixels and corresponding depths along back-projected rays. We compute a loss that enforces both accurate signed distances and consistent gradients, and update the MLP parameters via back-propagation. This process yields continuous neural SDFs suitable for downstream robotics tasks. Details on the network architecture and sampling strategy are in section IV-A2, and our loss function is described in section IV-A4.
@@
Figure 3: Block diagram illustrating the pipeline of safe reactive navigation using SDFs from RGB-D based Neural Barrier Functions.
@@ IV-A2 Neural Architecture
The neural SDF (figure 2) is defined by a parameterized MLP, $h_{\text{sdf}}(\xi;\theta)$, where $\xi\in\mathbb{R}^{3}$ denotes the input coordinates, $H$ denotes the hidden layers of MLP and $\theta=\{\mathbf{W}^{(i)},\mathbf{c}^{(i)}\}_{i=1}^{H}\cup\{\mathbf{W},\mathbf{c}\}$ represents the learnable weight matrices and biases. For brevity, we will write $h_{\text{sdf}}(\xi)$ instead of $h_{\text{sdf}}(\xi;\theta)$, recognizing that $\theta$ is implicitly updated during training. It comprises four hidden layers, each with 256 nodes, utilizing the Softplus activation with a high $\beta$ in intermediate layers to ensure smooth gradient updates while preserving non-linearity. The final output layer remains linear to maintain the integrity of the SDF values. To enhance the network’s ability to capture high-frequency details, we apply positional embedding (PE) to the input coordinates before passing them into the network. The PE, formally defined in (10), is also concatenated to the third layer. This “off-axis” embedding, inspired by Mip-NeRF 360 [25], employs periodic activation functions to project the input into a higher-dimensional space, thereby improving the reconstruction of fine-grained details.
[…]
IV-A3 Sampling Strategy
The sampling strategy follows the iMAP strategy [26] at the frame level and the NeRF[17] at the point level. To optimize computational efficiency and avoid redundancy, the system does not process all incoming frames but instead selects a curated subset. Similar to [20], keyframes are maintained through active sampling to mitigate memory loss. Initially, the first frame is always stored, while subsequent frames are selected based on their information gain, computed using a static version of the model. Each iteration consists of five frames: the two most recent frames and three keyframes chosen from the stored set, weighted by a normalized distribution of their cumulative losses. This selection cycle is repeated ten times before incorporating newer frames. Within each selected frame containing a camera pose $T$ ($T\in\mathbb{R}^{4\times 4}$) and an associated depth image $D$ ($D\in\mathbb{R}^{h\times w}$). The system samples points within the camera’s viewing frustum. First, 200 pixel coordinates $[u,v]$ are randomly selected from the depth image. Using the camera intrinsic matrix $K$ ($K\in\mathbb{R}^{3\times 3}$), viewing directions are computed and transformed into world coordinates as, $r=TK^{-1}[u,v,1]^{\top}$. For each ray $r$, we sample $N+M+1$ depth values $d_{i}$, including $N$ stratified samples from $[d_{\text{min}},D[u,v]+\hat{\delta}]$ and $M$ Gaussian-distributed samples from $\mathcal{N}(D[u,v],\sigma^{2})$ to improve supervision near the surface. Additionally, the surface sample from the depth map, $d=D[u,v]$, is always included. These samples define the 3D points along each ray as: $\mathbf{\xi_{i}}=d_{i}\cdot r$. The collected points from all selected frames form the training batch. In our experiments, we set $N=19$, $M=8$, $d_{\text{min}}=7\,\text{cm}$, $\hat{\delta}=10\,\text{cm}$, and $\sigma=10\,\text{cm}$.
IV-A4 Loss Function
@@
Here, $b$ represents the estimated ground truth SDF value. The loss is zero if the prediction is within the bound $b$, linear when exceeding $b$, and exponential for negative values. Since computing exact signed distances for all surface points is computationally expensive, we employ the batch distance method [20] to approximate the ground-truth SDF efficiently while balancing performance and computational overhead.
Near the surface, the loss should be more stringent. The near-surface loss, $L_{\text{near\_surf}}$, directly supervises the SDF prediction within these confines and is formulated using a Huber loss[22]:
@@
$\begin{cases}\frac{1}{2}(h_{\text{sdf}}(\xi)-b)^{2},&|h_{\text{sdf}}(\xi)-b|\leq\delta,\\ \delta(|h_{\text{sdf}}(\xi)-b|-\frac{1}{2}\delta),&|h_{\text{sdf}}(\xi)-b|>\delta.\end{cases}$
Previous works [20] primarily use L1 loss, but our experiments demonstrate that Huber loss mitigates the impact of noisy depth values, improving the quality of neural SDF reconstruction. A qualitative analysis comparing Huber and L1 loss under real-world noisy depth settings is presented in Section V-C. The combined SDF loss is given by:
@@
$\begin{cases}\lambda_{\text{surf}}L_{\text{near\_surf}}(h_{\text{sdf}}(\xi),b,\delta),&|D[u,v]-s|<\epsilon,\\ L_{\text{free\_space}},&\text{otherwise}.\end{cases}$
The second component, the gradient loss, refines the SDF prediction by supervising its gradient. This is crucial for ensuring smooth and accurate gradient behavior. To enforce gradient consistency, the gradient loss penalizes the cosine distance between the predicted gradient and the approximated gradient $\mathbf{g}$:
$L_{\text{grad}}(h_{\text{sdf}}(\xi),\mathbf{g})=1-\frac{\nabla_{x}h_{\text{sdf}}(\xi)\cdot\mathbf{g}}{\|\nabla_{x}h_{\text{sdf}}(\xi)\|\|\mathbf{g}\|}.$ (14)
@@
$\begin{cases}0,&\|\nabla_{x}h_{\text{sdf}}(\xi)\|-1<\alpha^{\prime},\\ \|\|\nabla_{x}h_{\text{sdf}}(\xi)\|-1\|,&\text{otherwise}.\end{cases}$
@@
where $\lambda_{\text{grad}}$ and $\lambda_{\text{eik}}$ control the relative influence of gradient and Eikonal constraints. For training, we use the following hyperparameters: $\beta=100$, $\alpha^{\prime}$ = 0.1, learning rate = $0.0013$, $b=30$ cm, and weight decay = $0.012$, $\lambda_{\text{surf}}$ = 5.3, $\lambda_{\text{grad}}$ = 0.02, $\lambda_{\text{eik}}$ = 0.27.
IV-A5 Continuity & Differentiability of Neural SDF
As noted in Section IV-A2, our neural SDF $h_{\text{sdf}}(\xi)$ first embeds the input $\xi\in\mathbb{R}^{3}$ into a higher-dimensional space via PE (10), which employs smooth (infinitely differentiable) $\sin$ and $\cos$ functions [25]. We then feed $\gamma(\xi)$ into an MLP with $H$ hidden layers. At each layer $i$, an affine map $\mathbf{z}_{i}=\mathbf{W}^{(i)}\mathbf{q}_{i-1}+\mathbf{c}^{(i)}$ is followed by a $\operatorname{Softplus}$ activation $\mathbf{q}_{i}=\operatorname{Softplus}\bigl{(}\mathbf{z}_{i};\beta\bigr{)}$, where $\operatorname{Softplus}(z;\beta)=\frac{1}{\beta}\ln\bigl{(}1+e^{\beta z}\bigr{)}$ [27]. We set the initial layer to $\mathbf{q}_{0}=\gamma(\xi)$. The final output layer remains linear to preserve the signed-distance property:

### v3

@@ Abstract
Autonomous safe navigation in unstructured and novel environments poses significant challenges, especially when environment information can only be provided through low-cost vision sensors. Although safe reactive approaches have been proposed to ensure robot safety in complex environments, many base their theory off the assumption that the robot has prior knowledge on obstacle locations and geometries. In this paper, we present a real-time, vision-based framework that constructs continuous, first-order differentiable Signed Distance Fields (SDFs) of unknown environments entirely online, without any pre-training, and is fully compatible with established SDF-based reactive controllers. To achieve robust performance under practical sensing conditions, our approach explicitly accounts for noise in affordable RGB-D cameras, refining the neural SDF representation online for smoother geometry and stable gradient estimates. We validate the proposed method in simulation and real-world experiments using a Fetch robot. Videos and supplementary material are available at https://satyajeetburla.github.io/rnbf/.
@@ I INTRODUCTION
Safety in robotics is a critical challenge, especially when operating in unstructured environments where obstacle states and geometries are unavailable [yang2018grand]. In recent years, Control Barrier Functions (CBFs) were widely studied due to their ability to provide safety guarantees for any control affine systems and adopted in various applications including vehicle cruise control [cbfcruise], indoor navigation [xue2025minimacollisionscombiningmodulation], quadrotor control [cbfquadrotor], legged robot locomotion [grandia2021multi], etc. All the aforementioned works assume that the robot has prior knowledge of obstacle states and geometries, or even a predefined map. However, for robot navigation in unknown environments, such assumptions are unrealistic and the robot can only access noisy environment information from its onboard sensors.
CBFs’s performance in enforcing robot safety relies not only on the robot’s ability to collect accurate nearby obstacle state information but also on how well the collected information is processed to meet the requirement of valid CBFs. Therefore, when environment descriptions are unavailable or noisy, the safety guarantee of robot navigation is jeopardized. Building valid CBFs adaptively using noisy data collected from robots’ on-board sensors in real time remains a challenge. Many recent works use LiDAR sensors to construct CBFs and ensure robot safety in unknown environments [zhang2024online, dawson2022learning, long2021learning, harms2024neural]. In this work, we instead demonstrate that safe reactive control can be achieved with RGB-D camera alone, without relying on additional sensing modalities such as LiDAR.
@@
Seldom works have investigated perception-based CBFs using RGB-D camera inputs. Existing vision-based CBF pipelines utilize machine-learning techniques to incorporate robot observations into CBF formulations. For example, [abdi2023safe] leverages a conditional Generative Adversarial Network to approximate CBFs given RGB-D camera inputs; [tong2023enforcing] predicts future RGB-D observations with Neural Radiance Fields (NeRFs) and feeds them into discrete-time CBFs. These methods require significant offline training and their performance can be affected by the quality of their training dataset. Additionally, using CBFs in discrete-time settings may weaken their safety guarantees when the update frequency is not fast enough. Other works, like [de2024point], approximate mathematically CBFs and their gradients from point cloud data collected through LiDARs or depth cameras.
Among the methods listed above, it is noteworthy that none has explicitly considered the effects of sensory noises on perception-based CBF. Further, most existing vision-based CBFs have made approximations to standard CBFs to accommodate robot observation formats. Assumptions that CBFs are founded on, such as the continuity and first-order differentiability of the barrier function are typically relaxed, weakening their safety guarantees. Additionally, altering standard CBFs in vision-based safe navigation hinders the adaptation of well-developed none-sensor-based safe reactive control algorithms into vision-based robot platforms. Many safe reactive controllers, including Control Barrier Function Quadratic Programs (CBF-QP), are developed assuming the system’s access to SDFs and their gradients of the obstacles near the robot [zucker2013chomp],[cbfcruise],[cbfstar],[LukesDS], [onManifoldMod] and [xue2025minimacollisionscombiningmodulation]. In robot safe navigation, SDFs refer to the orthogonal distance of a given point to the boundary of an unsafe set.
Therefore, in this paper, we offer an online barrier function generation framework capable of smoothing sensor noises and constructing continuous and first-order differentiable neural SDFs in real-time given RGB-D camera inputs without pre-training for any unknown environments. With the advancements presented by NeRF[mildenhall2021nerf], the quality and efficiency of environmental reconstruction have significantly improved. However, limited work explored generating SDFs using NeRF, most of which were designed for an offline setting [liu2022regularized, tang2023rgb]. A notable exception is iSDF[ortiz2022isdf], which introduced a method for real-time SDF generation. Recently, [bukhari2025differentiable] proposed a dual-level SDF approach combining pretrained object SDFs with a scene-level iSDF for indoor navigation. Their method focused on trajectory optimization and relies on offline components, without incorporating any safety-filtering framework. In contrast, our vision pipeline generates real-time, continuous, and differentiable SDFs, specifically designed for constructing continuous CBFs for safe robot navigation. Instead of altering CBFs to accommodate vision observations, as in prior works, we establish continuous and first-order differentiable SDFs online, which enables our methods to be readily adaptable to SDF-based safe reactive controllers such as standard CBF-QPs [ames2019control, xue2025no] and dynamical system based modulation [fourie2024manifold, koptev2024reactive]. Our approach is validated in both gazebo simulations and real-life hardware experiments using the Fetch robot in static and quasi-static obstacle environments.
Our key contributions are as follows:
  • We introduce an approach to generate real-time (5-15 Hz) SDFs from raw RGB-D data using an online neural representation compatible with established SDF-based reactive controllers.
  • To account for depth noise inherent in affordable sensors like RGB-D cameras, we mitigate its impact on neural SDF representations adopting the huber loss [huber1992robust].
@@ II Problem Formulation
Given a continuously differentiable function $\mathbb{R}^{n}\rightarrow\mathbb{R}, hh$ is a barrier function if the safe set $C$ (outside the obstacle), the boundary set $\partial C$ (on the boundary of the obstacle), and the unsafe set $\neg C$ (inside the obstacles) of the system can be defined as in (2), (3), (4[ames2019control].
@@
Figure 2: Block diagram of the proposed RNBF-based safe reactive navigation pipeline. The SDF Generation module (5–15 Hz) constructs a continuous, differentiable signed distance field from real-time RGB-D input, while the Reactive Controller ($\geq 20$ Hz) ensures collision-free navigation using a CBF-QP. The black thick arrow highlights the main coupling step between the perception and control pipelines. Red outlines highlight components in the SDF generation stage that distinguish the RNBF pipeline from the original iSDF pipeline.
@@ III-A Control Barrier Function
C $C\text{ is set invariant}\iff\dot{h}(x)\geq 0\;\forall x\in\partial$ C\\ (5)
Control Barrier Functions are designed by extending Nagumo set invariance theorem to a “control” version[ames2019control], where the condition $\forall x\in\partial C$ is rewritten mathematically using an extended $\mathcal{K}_{\infty}$ function $\alpha$,
@@ C $C\text{ is set invariant}\iff\dot{h}(x)\geq 0\;\forall x\in\partial$ C\\ (5)
An extended $\mathcal{K}_{\infty}$ function is a function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ that is strictly increasing and with $\alpha(0)=0$ ; that is, extended $\mathcal{K}_{\infty}$ functions are defined on the entire real line $(-\infty,\infty)$.
@@
[…]
$L_{f}h(x)+L_{g}h(x)u\geq-\alpha(h(x))$ (8)
@@ III-B Signed Distance Function (SDF)
When representing three-dimensional geometry, a signed distance field $h_{\text{sdf}}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ is a scalar field that maps a 3D coordinate $\xi=[p_{x},p_{y},p_{z}]^{\top}$ to the scalar signed distance value [ortiz2022isdf]. The surface $S$ is the zero level-set of the field:
@@
III-C iSDF: Real-time SDF Reconstruction
iSDF[ortiz2022isdf] proposed a continual learning system for real-time estimates a SDF from a stream of depth images with known camera poses. The SDF is represented by a multilayer perceptron (MLP) mapping a 3D location to its signed distance. At each iteration, the system selects a small set of depth frames via active keyframe replay to mitigate catastrophic forgetting. For each selected frame, pixels and depths are sampled along back-projected rays to form a batch of 3D query points. The model is evaluated at these points and trained with a self-supervised objective; parameters are updated by backpropagation. The work reported real-time performance and evaluates the method on real and synthetic indoor datasets[ortiz2022isdf].
@@ IV-A1 Core Methodology
Our real-time SDF pipeline (Figure 2, “SDF Generation”) follows the iSDF[ortiz2022isdf] template but differs primarily in two major components that are critical for safe robot navigation: (i) an explicit floor segmentation stage (ii) a robust near–surface loss (Sec. IV-A3) that replaces the $L_{1}$ term used in iSDF to attenuate RGB-D depth noise.
Given a sequence of posed RGB-D images, Our pipeline first segments the floor and masks the corresponding depth pixels to prevent the large planar ground from being treated as an obstacle, following standard RGB-D scene understanding practice [gupta2013perceptual]. The segmentation module is modular, and any modern method (e.g., SAM [kirillov2023segment]) can be used. In our implementation, for simplicity we use a fast color-based segmenter on the RGB image and apply the resulting mask to the depth image. At the heart of our pipeline also lies a MLP, which transforms a three-dimensional point $\xi$ into its corresponding signed distance. The network (Sec. IV-A2) is randomly initialized and refined online from the incoming masked depth stream. We also maintain a small keyframe set, select a few frames per update via active replay, and sample query points along back-projected rays using the sampling strategy as in iSDF[ortiz2022isdf]. The neural SDF $h_{\text{sdf}}(\xi)$ is queried at these sampled points and optimized with the objective in section IV-A3, yielding a continuous SDF and gradients available online (5–15 Hz) to the reactive safety controller (running at $\geq 20$ Hz) for safe robot navigation.
@@ IV-A2 Neural Architecture
The neural SDF (Figure 3) is defined by a parameterized MLP, $h_{\text{sdf}}(\xi;\theta)$, where $\xi\in\mathbb{R}^{3}$ denotes the input coordinates, $H$ denotes the hidden layers of MLP and $\theta=\{\mathbf{W}^{(i)},\mathbf{c}^{(i)}\}_{i=1}^{H}\cup\{\mathbf{W},\mathbf{c}\}$ represents the learnable weight matrices and biases. For brevity, we will write $h_{\text{sdf}}(\xi)$ instead of $h_{\text{sdf}}(\xi;\theta)$, recognizing that $\theta$ is implicitly updated during training. It comprises four hidden layers, each with 256 nodes, utilizing the Softplus activation with a high $\beta$ in intermediate layers to ensure smooth gradient updates while preserving non-linearity. The final output layer remains linear to maintain the integrity of the SDF values. Finally, inspired by Mip-NeRF 360 [barron2022mip] & iSDF[ortiz2022isdf], we also apply positional embeddings(PE) to the input coordinates $\xi$ and concatenate the embedding at an intermediate layer (the third layer) to better capture high-frequency detail.
IV-A3 Loss Function
@@
Here, $b$ represents the estimated ground truth SDF value. Since computing exact signed distances for all surface points is computationally expensive, we employ the batch distance method [ortiz2022isdf] to approximate the ground-truth SDF efficiently while balancing performance and computational overhead.
Near the surface, the loss should be more stringent. The near-surface loss, $L_{\text{near\_surf}}$, directly supervises the SDF prediction within these confines and is formulated using a huber loss[huber1992robust]:
@@
[…]
Previous works like iSDF[ortiz2022isdf] primarily use L1 loss, but our experiments demonstrate that huber loss mitigates the impact of noisy depth values, improving the quality of neural SDF reconstruction. A qualitative analysis comparing huber and L1 loss under real-world noisy depth settings is presented in section V-C. The combined SDF loss is given by:
@@
[…]
The second component, the gradient loss, refines the SDF prediction by supervising its gradient and ensure smooth, accurate gradient. To enforce gradient consistency, the gradient loss penalizes the cosine distance between the predicted gradient and the approximated gradient $\mathbf{g}$:
$L_{\text{grad}}(h_{\text{sdf}}(\xi),\mathbf{g})=1-\frac{\nabla_{\xi}h_{\text{sdf}}(\xi)\cdot\mathbf{g}}{\|\nabla_{\xi}h_{\text{sdf}}(\xi)\|\|\mathbf{g}\|}.$ (13)
@@
[…]
@@
where $\lambda_{\text{grad}}$ and $\lambda_{\text{eik}}$ control the relative influence of gradient and Eikonal constraints.
IV-A4 Continuity & Differentiability of Neural SDF
As noted in section IV-A2, our neural SDF $h_{\text{sdf}}(\xi)$ first embeds the input $\xi\in\mathbb{R}^{3}$ into a higher-dimensional space via PE, which employs smooth (infinitely differentiable) $\sin$ and $\cos$ functions [barron2022mip]. We then feed $\gamma(\xi)$ into an MLP with $H$ hidden layers. At each layer $i$, an affine map $\mathbf{z}_{i}=\mathbf{W}^{(i)}\mathbf{q}_{i-1}+\mathbf{c}^{(i)}$ is followed by a $\operatorname{Softplus}$ activation $\mathbf{q}_{i}=\operatorname{Softplus}\bigl(\mathbf{z}_{i};\beta\bigr)$, where $\operatorname{Softplus}(z;\beta)=\frac{1}{\beta}\ln\bigl(1+e^{\beta z}\bigr)$ [glorot2011deep]. We set the initial layer to $\mathbf{q}_{0}=\gamma(\xi)$. The final output layer remains linear to preserve the signed-distance property: